OpenStudy (anonymous):

Your piggy bank has a total of 43 coins in it; some are dimes and some are nickels. If you have a total of $3.60, how many dimes do you have?

OpenStudy (anonymous):

Let no of dimes= d let no of nickels= n total coins in piggy bank ( n+d)= 43...(equation 1) a dime = 10 cent a nickel = 5 cent equation 2: 5n+10d=360 equation 1: n+d=43 two equations and two variables now we can calculate the "d and n". after calculation dimes(d)=29 nickels(n)=14